OpenStudy (anonymous):

when we add 2+3+7-10=

Parth (parthkohli):

We get 42.

Parth (parthkohli):

jk, we get 5 + 7 - 10

OpenStudy (unklerhaukus):

4/2

OpenStudy (lgbasallote):

= 2 + 3 + 7 - 10 = 5 + 7 - 10 = 12 - 10 = 2